GAS M&O meeting 26 March 2020
Present: Beatrice, Roberto, Andrea, Patrick, Kacper, LPh, Fred, Michal, Mara, Gianluca, Sabino, Fabio, Federico, Leonardo
Indico: https://indico.cern.ch/event/902256/
 
CHECK Gas Systems
 
 
GENERAL
 
	- Check alarm levels for PGS —> All gas system responsible
 
	- To prepare a list of sensors not working or missing in PGS —> All gas system responsible
 
	- Check with David for aging test of new sensors used in PGS —> Beatrice
 
	- CMS DT pump: to plan intervention to change the membrane —> Andrea + Kacper
 
	- CMS CSC batteries —> Andrea to keep an eye
 
	- Discussion about new IR module
	
		- The pressure regulator will be only reading in WinCC-OA but it will be possible to regulate it on place with the black display
 
		- Addition of small pump in IR module: still under discussion
 
	
	 
	- Excel file for alarms prepared weekly by Mara —> Experiment technical responsible to fill
 
	- Excel file for ALICE and ATLAS —> Gianluca to prepare with LPh and Kacper
 
 
 
 
 
REMINDER for when we will be back
 
ALICE
CPV
	- We have 3 lines input that are ok, and 3 lines output: in one there is not the return. LPh contacted Yuri
 
TPC
	- In the purifier the small thermic box triggered (also last year it happened) —> Patrick/LPh to check
 
TRD
	- Recuperation: missing two valves but there are present in the software
 
	- Purifier: to review if keep the safety valve or the rupture disk
 
	- Purifier: safety valve at the input where we have Ar or Ar/H2 at high pressure —> LPh to check with safety
 
 
ATLAS
MDT
	- Check overpressure in MDT chamber when there is a stop-start of the system
 
	- Only one convertisseur, one pump, etc from the system —> Patrick/Kacper to fix
 
	- Found a problem of grounding on pressure sensors: pressure sensors changed. Now to verify if the grounding problem is still there
 
	- Problem in the 2 vaisala of humidifier: several spikes present. Not clear why. Before the software upgrade they were not present
 
 
CMS
RPC
	- Michal changed the mother board in rack 67: now flow is 170 l/h (elog #2279) —> Andrea to check with new voegtlin red
 
DT
	- Change one motherboard in rack 61 but flow didn’t change (elog #2280) —> to try to recalibrate one set of flowcells
 
	- Difficult to re-start the gas system after power-cut —> to do some tests
 
	- Test for power cut and re-start for the exhaust line —> to do other tests
 
GEM
	- Andrea prepared a drawing for ME0 and check for space in UX —> Andrea to send an email to Michele
 
CSC
	- On-going regulation of PID for the exhaust —> Beatrice
 
	- Warning to set-up —> Beatrice/Michal/Roberto
 
 
LHCb
 
MWPC
	- Problem in the GC analysis of the mixer: not visible the correct gas mixture with both LF and HF. Probably an hardware problem (connection, valve, etc) to find —> Mara
 
	- Leak in Mixer and Exhaust. Only in Exhaust found but not repaired. Other not yet founded —> Patrick
 
RICH2
	- To inject Air —> Patrick to check with C. Frei
 
	- To add line for analysis “supply to detector”k
 
SciFi
	- Problem on 3-ways valve —> Patrick
 
PGS
	- Interlock of Ar/H2 does not work —> Michal/Patrick to check 
 
 
GENERAL
 
	- Alarms/Errors in power supplies for all gas systems —> Patrick
 
             
            
                
                    There are minutes attached to this event.
                    
Show them.